The ages of Edna, Ellie, and Elsa are consecutive integers. The sum of the ages is 111. what are their ages?
Answer by waynest(281)

edna = x
ellie = x + 1
elsa = x + 2
x + (x + 1) + (x + 2) = 111
3x + 3 = 111
3x + 3 - 3 = 111 - 3
3x = 108
3x/3 = 208/3
x = 36
check:
36 + (36 + 1) + (36 + 2) = 111
36 + 37 + 38 = 111
111 = 111